OpenStudy (anonymous):

A local snail population grows according to the function g(x)=200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then explain to Iris how the key features of this local snail population compares to the key features of the invasive population. @Directrix @sammixboo @Compassionate @eliassaab @ganeshie8 @Nnesha @SolomonZelman

OpenStudy (solomonzelman):

Apologize for a late reply. \(\large\color{slate}{ g(x)=200(1.03)^{2x} }\) \(\large\color{slate}{ g(x)=200\left(\begin{matrix} ~(1.03)^{2} ~ \\ \end{matrix}\right)^x }\)

OpenStudy (solomonzelman):

then simplify inside the parenthesis.

OpenStudy (anonymous):

is that all

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

it is in terms of an exponent of a single x.

OpenStudy (solomonzelman):

so you made a function with an exponent of x, but didn't change it's value:)